OpenStudy (anonymous):

Dennis and Darla have a 15/6 balloon mortgage for $178,800 with a rate of 3.95%. How much will they pay in interest over the life of the loan?

OpenStudy (anonymous):

is the 15/6 a ratio?

OpenStudy (anonymous):

refers to the terms of the mortgage

OpenStudy (anonymous):

15 means that the beginning fixed monthly payments are going to be calculated as if the mortgage would be paid off in 15 years

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

and the 6 means?

OpenStudy (tkhunny):

You calculate the mortgage payment based on 15 years. You pay this 15=year amount for six years only. The balance is due at the end of six years. This would be the "balloon" part.

OpenStudy (anonymous):

@Hope_nicole The 6 means that fixed payments will be made for 6 years. After the 6 years, the balloon payment is due.

OpenStudy (anonymous):

okay, so the life of the loan is 6 years..a mortgage(aka starting value) of 178800 and a rate of 3.95, correct?

OpenStudy (tkhunny):

The life of the loan is six years, but the payments during those six years are calculated as if the life of the loan were 15 years.

OpenStudy (anonymous):

@Hope_nicole correct

OpenStudy (anonymous):

@sunil186 Joanne and Bill are financing $305,500 to purchase a house. They obtained a 20/6 balloon mortgage at 5.75%. What will their balloon payment be?

OpenStudy (tkhunny):

I don't see any progress, here. Can you calculate the payment for the 20-year mortgage?

OpenStudy (anonymous):

@tkhunny i have no idea how to do this. I'm soooooo confused

OpenStudy (tkhunny):

Seriously? You must step up your game a little. financing $305,500 20/6 balloon mortgage 5.75% I = 0.0575 j = i/12 = 0.004792 v = 1/(1+j) = 0.995231 20 year mortgage \(305500 = Pmt*(v + v^2 + ... + v^240) = \dfrac{v - v^{241}}{1-v}\) Solve that to find the Payment for the 20-year mortgage. Pay this for six years. \(Balloon Payment = Pmt*\dfrac{v - v^{169}}{1-v}\) Really, they only get trickier, not easier. You must be very clear on all of this.

OpenStudy (anonymous):

@tkhunny I got $36,132.26 ...

OpenStudy (tkhunny):

You will have to provide MUCH more information. That is not very close and I have no idea why. You didn't share anything. Reality Check: You honestly believe that 39132.26 / 305500 = 13% only 13% of a 20 year loan remains after 6 years? Try more like 80% Things have to make sense.